OpenStudy (anonymous):

CAN SOMEONE HELP ME FIND THE LCD THAT WILL ISOLATE THE VARIABLE IN THE EQUATION , X/2+2/3=3/4 THANKS!

OpenStudy (anonymous):

3=1?

Directrix (directrix):

X/2 +2/3 = 1/4 The LCD will need a 2, a 3, and another 2. The first 2 comes from x/2. The three comes from 2/3. Where you see the 4 in 1/4, you might think 4 is a factor of the LCD because 4 = 2*2. But, there is already one 2 from the first term. Only one more 2 is needed as a factor. So, multiply 2 times 3 times 2 to get the LCD. @jenny1

OpenStudy (anonymous):

No that wasnt it.

OpenStudy (anonymous):

ok see what u need to do is

OpenStudy (anonymous):

multiplay all D

OpenStudy (anonymous):

the least common multiple of 3 and 4 is 3 times 4 which is 12

Directrix (directrix):

@jenny1 >>No that wasnt it. Why do you say that?

OpenStudy (anonymous):

\[\frac{x}{2}+\frac{2}{3}=\frac{3}{4}\] \[12\left (\frac{x}{2}+\frac{2}{3}\right)=12\times \frac{3}{4}\] \[6x+8=9\]solve for \(x\)